OpenStudy (anonymous):

If there are 5 sciene and 3 math teachers, what is the probability that out of three teachers, one will be a math theacher

OpenStudy (anonymous):

\[\frac{\dbinom{5}{2}\dbinom{3}{1}}{\dbinom{8}{3}}\]

OpenStudy (anonymous):

finally \[\frac{10\times 3}{56}\]